OpenStudy (anonymous):

Using the replacement set, find the solutions for the equation. y = 0.5x + 10 Replacement Set: {(-2, 9.5), (0, 10),(3, 11.5),(4, 13)} A. {(-2, 9.5), (0, 10),(3, 11.5),(4, 13)} B. {(-2, 9.5), (0, 10),(3, 11.5)} C. {(0, 10),(3, 11.5),(4, 13)} D. { (0, 10),(3, 11.5)}

OpenStudy (jdoe0001):

all it means is set "y" and "x" according to the pair given, like say \(\bf \color{blue}{y} = 0.5\color{red}{x} + 10 Replacement Set: {(\color{red}{-2}, \color{blue}{9.5}), (\color{red}{0}, \color{blue}{10}),(\color{red}{3}, \color{blue}{11.5}),(\color{red}{4}, {13})}\) so lemme get you the 1st pair for the solution say \(\bf y = 0.5x + 10\qquad \qquad (-2, 9.5)\qquad x = -2\qquad y = 9.5\\ \quad \\ (9.5) = 0.5(-2)+10\implies 9.5 = -1+10\implies 9.5 = 9\quad ?\) so.... notice we're using an EQUATION, so the left side and right side must EQUATE if the EQUATion is to be valid or TRUE so, is 9.5 really equals 9? no, thus is NOT A SOLUTION try the other given pairs, see which one gives you a valid EQUATion